OpenStudy (anonymous):

e^2x-7e^x+10=0 Show all steps

hero (hero):

Hint: (e^x)^2 - 7e^x + 10 = 0 Let e^x = u u^2 - 7u + 10 = 0

OpenStudy (anonymous):

How did you turn it into a u?

OpenStudy (anonymous):

That hint didn't help

OpenStudy (anonymous):

u^2 - 7u + 10 = 0 is u = 2,5 But the answer for e^2x-7e^x+10=0 is x = -1.0074 and x = 0.83

hero (hero):

I give obvious hints and students are still clueless.

OpenStudy (anonymous):

let \[e^x=u\] so\[(e^x)^2-7e^x+10\] very obvious how would you substitute for u

hero (hero):

That's because you did it wrong bro @YesterdayiSaidTomorow

OpenStudy (anonymous):

So... I raise the answers of u^2 - 7u + 10 = 0 to the second power? I don't get what you're saying

hero (hero):

You're supposed to factor u^2 -7u + 10 to get (u - 5)(u - 2) = 10 Then re-write as: u - 5 = 0 u - 2 = 0 Then RE-SUBSTITUTE e^x back in: e^x - 5 = 0 e^x - 2 = 0 Now continue solving for x:

OpenStudy (anonymous):

ooooo

OpenStudy (anonymous):

can you finish it to ensure I get the right answer?

hero (hero):

Nope. You already have the answers. Figure it out bro.

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

fiesty

OpenStudy (anonymous):

Lol, but I got log(5)

OpenStudy (anonymous):

and log(2)

OpenStudy (anonymous):

ln

OpenStudy (anonymous):

???

OpenStudy (anonymous):

you dont use log for e,log is for base10 ln is for base e

OpenStudy (anonymous):

Ugh I don't know how to do this at all

OpenStudy (anonymous):

\[x=\ln(5) ,x=\ln(2)\]

OpenStudy (anonymous):

but that doesn't match http://www.wolframalpha.com/input/?i=e%5E2x-7e%5Ex%2B10%3D0

hero (hero):

Yeah, just use ln(5) and ln(2)

OpenStudy (anonymous):

that 1.6 and the other is -1.0074?

OpenStudy (anonymous):

Yahoo Answers said the same thing, so whatever and thank you!

hero (hero):

I get: ln(5) = 1.60944 ln(2) = 0.693147